OpenStudy (anonymous):

adding fractions with different denominators: solve 7/10+1/6= adding fractions with different denominators: solve 7/10+1/6= @Mathematics

OpenStudy (diyadiya):

52/60

OpenStudy (diyadiya):

13/15

OpenStudy (anonymous):

\[7/10+1/6=(42+10)/60=52/60\]

OpenStudy (anonymous):

thanks